OpenStudy (micheal1234):

which ordered pairs are solutions to the equation y = x – 4 a.(2,6) b.(6,2) c.(-2,6) d.(-6,2)

OpenStudy (dayakar):

b

OpenStudy (micheal1234):

i got more give me a sec

OpenStudy (dayakar):

if x= 6 y= x-4 =6-4 =2 (X,Y) =(6,2)

OpenStudy (dayakar):

IS IT CLEAR

OpenStudy (micheal1234):

same thing y=5x

OpenStudy (dayakar):

WHAT ARE THE OPTIONS

OpenStudy (micheal1234):

a. (-2,10) b.(-5,15) c.(-3,25) d.(-2,10)

OpenStudy (dayakar):

all are wrong

OpenStudy (micheal1234):

sorry b and c are b(-5,25) C(-3,,15)

OpenStudy (dayakar):

these also wrong

OpenStudy (micheal1234):

would it be (-2,-10)

OpenStudy (dayakar):

right

OpenStudy (micheal1234):

my typing skill are not good to day

OpenStudy (dayakar):

if x=-2 y=5x y= 5(-2) = -10 (-2,-5)

OpenStudy (dayakar):

no problem ,is it ok

OpenStudy (micheal1234):

y=-5x+4

OpenStudy (micheal1234):

a.(2,14) b.(5,-29) C.(3,-19) d. (5,-21)

OpenStudy (dayakar):

d

OpenStudy (dayakar):

over

OpenStudy (micheal1234):

solve the equation 6.4=4x

OpenStudy (dayakar):

x=6.4/4 x=1.6

OpenStudy (micheal1234):

thanks

OpenStudy (dayakar):

it's ok